OpenStudy (anonymous):

don't understand how y=2sin(3x + pie)+1 becomes 2sin(3(x-(-pie/3)+1. The formula is A sin (B (x-C) +D.

geerky42 (geerky42):

If you simplify \(2\sin\left(3\left(x-\left(-\dfrac{\pi}{3}\right)\right)\right)+1\), you would get \(2\sin(3x + \pi)+1\)